OpenStudy (anonymous):

What are the first four terms of the sequence defined by the following equation? an = 5·3n - 1 5 , 15 , 225 , 3,375 ... 15 , 45 , 135 , 405 ... 5 , 15 , 45 , 135 ... 15 , 225 , 3,375 , 50,625 ...

OpenStudy (campbell_st):

well subtitute n = 1 and see what you get \[a{1} = 5 \times 3^{1 - 1} = 5 \times 3^0 = 5\] repeat the process for n = 2, n = 3 and n = 4

OpenStudy (mathmale):

Your equation is a bit hard to read and interpret. I suspect you meant\[a _{n}=5*3^{n-1}\]Is that the correct interpretation?

OpenStudy (anonymous):

Yes